# Descent spectral sequences through synthetic spectra

### v1

@@ 3.3   Comparing the DSS and the ANSS
    Using the notation of (3.12), we have an isomorphism of graded abelian groups [^8] As usual, we set $\omega^{n/2}=0$ for odd $n$.
@@
Using part (2) of Pr. 3.20, we see that assumption (b) is equivalent to the analogous vanishing condition for the $\mathrm{E}_{2}$-page, so (b) certainly implies (a). To prove the converse, suppose that the group $\mathrm{H}^{s}(\mathsf{Y},\omega_{\mathsf{Y}}^{\otimes t})$ does not vanish for certain $s,t$. By part (2) of Pr. 3.20, this group appears on the $\mathrm{E}_{2}$-page of the SDSS in tridegree $(2t-s,\,s,\,s)$. We claim that by part (3) of Pr. 3.20, this tridegree cannot support nor be the target of any differentials. Indeed, it cannot be the target of a $d_{r}$-differential for $r\geq 2$, for such a differential only hits elements that are $\tau^{r-1}$-divisible on the $\mathrm{E}_{2}$-page, and this tridegree is not $\tau$-divisible. Note that by (2), the $\mathrm{E}_{2}$-page of the SDSS is concentrated in tridegrees $(x,y,z)$ where $y\leq z$. This means that the tridegree $(2t-s,\,s,\,s)$ cannot support any $d_{r}$-differentials for $r\geq 2$, as such a differential has tridegree $(-1,1,r)$. We therefore find that this tridegree gives a nonzero contribution to $\nu\mathrm{MU}_{2t-s,\,s}(\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X}))$. If $s>n$, this means that $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$ is not $(-n)$-connective. ∎
@@
This identification of spectral sequences, and other similar statements, have appeared in the literature before:
@@ 4.1   Homotopy fixed-point spectral sequences
that admits a canonical even-periodic refinement given by $\operatorname{Spec}(E)/G$; see [MM15, Pr.2.15]. It is well known that the homotopy fixed-point spectral sequence (HFPSS) for $E$ can be naturally identified with the DSS for the global sections of this even-periodic refinement, which we write as $(\mathsf{X},\mathcal{O}^{\mathrm{top}}_{\mathsf{X}})$ for a moment. Indeed, the cobar resolution computing the homotopy fixed points $E^{hG}$ takes the form of the cosimplicial object
@@
Let $G$ be a finite group and let $E\in\mathrm{Fun}(BG,\mathrm{CAlg}(\mathrm{Sp}))$ whose underlying ring spectrum is Landweber exact. Then, using the notation above, there is a natural equivalence of cosimplicial $\mathbf{E}_{\infty}$-rings $E^{\bullet}\cong\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(\operatorname{Spec}(\pi_{0}E)^{\bullet})$. In particular, the HFPSS for $E^{hG}$ and the DSS for $\mathsf{X}$ are isomorphic, and the signature of the synthetic $\mathbf{E}_{\infty}$-ring $(\nu E)^{hG}$ is the HFPSS for $E$.
@@
Let $G$ be a finite group and let $E\in\mathrm{Fun}(BG,\mathrm{CAlg}(\mathrm{Sp}))$ whose underlying ring spectrum is Landweber exact. If, for every field-valued point $x$ of $\operatorname{Spec}\pi_{0}E$, the subgroup of $G$ stabilising $x$ acts faithfully on the formal group associated to $x$, then the canonical map
@@
is an equivalence of synthetic $\mathbf{E}_{\infty}$-rings. In particular, the signature of $(\nu E)^{hG}$ is the ANSS of $E^{hG}$.
@@
In general, regardless of whether or not $\operatorname{Spec}(\pi_{0}E)/G\to\mathsf{M}_{\mathrm{FG}}$ is affine, the signature of the synthetic spectrum $\nu(E)^{hG}$ is the HFPSS of $E$ by Cor. 4.2. This will not generally coincide with the ANSS of $E^{hG}$, making $\nu(E)^{hG}$ often preferable over $\nu(E^{hG})$ as a synthetic lift of $E^{hG}$. The unit of the synthetic $\mathbf{E}_{\infty}$-ring $\nu(E)^{hG}$ still provides a map of multiplicative spectral sequences
@@
suitable for detection arguments, for example. In this sense, the HFPSS for $E$ is a modified ANSS à la [CD24, Df.1.5]. From this point of view, the ANSS of $E^{hG}$ has no advantage over the HFPSS.
@@
In fact, the corresponding homotopy fixed-point spectral sequence does not coincide with the Adams–Novikov spectral sequence of $\mathrm{KU}^{hC_{2}}\cong\mathrm{KU}\otimes B{C_{2}}_{+}$. Indeed, the latter is complex-oriented as a $\mathrm{KU}$-module, and hence its ANSS is concentrated in filtration zero. On the other hand, the cohomology groups $\mathrm{H}^{s}(C_{2},\pi_{*}\mathrm{KU})$ are nontrivial for positive values of $s$, hence the homotopy fixed point spectral sequence is nontrivial in positive filtrations.
Since the ANSS of $\mathrm{KU}\otimes B{C_{2}}_{+}$ collapses on the zero-line, the ANSS provides no computational leverage; computing the $\mathrm{E}_{2}$-page already requires knowing $\pi_{*}(\mathrm{KU}\otimes B{C_{2}}_{+})$. The HFPSS, however, has an $\mathrm{E}_{2}$-page that is algebraically computable using only $\pi_{*}\mathrm{KU}$ as input.
@@ 4.2   Synthetic modular forms
By Pr. 2.14, the synthetic $\mathbf{E}_{\infty}$-rings $\mathrm{Smf}$ and $\mathrm{SMF}$ are synthetic lifts of $\mathrm{Tmf}$ and $\mathrm{TMF}$, and by Cor. 2.12, their signatures are the DSS for $\mathrm{Tmf}$ and $\mathrm{TMF}$.
@@
We will ignore the case of $\mathrm{tmf}$, as there is no clear way to define this object from the spectral algebraic geometry used here. The first main problem is that the stack $\mathsf{M}_{\mathrm{cubic}}$ associated to $\mathrm{tmf}$ is not Deligne–Mumford, as the automorphism group of the cuspidal cubic given by $y^{2}=x^{3}$ is not an étale group scheme. The second problem is that the map $\mathsf{M}_{\mathrm{cubic}}\to\mathsf{M}_{\mathrm{FG}}$ is not flat. To see this, we notice that if it were flat, then $\mathrm{MU}_{\ast}\mathrm{tmf}$ would be flat over $\pi_{\ast}\mathrm{MU}$, which we can see fails once we know that
@@
In a similar vein, $\operatorname{\nu}\mathrm{ko}$ is a preferred $\mathrm{MU}$-synthetic model for $\mathrm{ko}$ as the map $\mathsf{M}_{\mathrm{quad}}\to\mathsf{M}_{\mathrm{FG}}$ (where $\mathsf{M}_{\mathrm{quad}}$ is the stack associated to $\mathrm{ko}$; see [DFHH14, §9]) is also affine. In this case, the first two authors previously showed [CD24, Pr.3.14] by direct computation that there are equivalences
@@
which further confirm this suspicion that $\operatorname{\nu}\mathrm{ko}$ is the preferred synthetic lift of $\mathrm{ko}$. We repeat the warning from [CD24, Warn.3.16] that these equivalences do not hold in $\mathrm{Syn}_{\mathbf{F}_{2}}$.
@@
Variant 4.11.
It is also possible to alter Df. 4.8 to accommodate moduli stacks of elliptic curves with level structures. Indeed, for a congruence subgroup $\Gamma\subseteq\operatorname{SL}_{2}(\mathbf{Z})$, write $\mathsf{M}_{\Gamma}$ for the moduli stack of elliptic curves with $\Gamma$-level structure. The map $\mathsf{M}_{\Gamma}\to\mathsf{M}_{\mathrm{Ell}}$ is étale, and we define
@@
This is a synthetic $\mathbf{E}_{\infty}$-ring that lifts $\mathrm{TMF}_{\Gamma}$ and whose signature is the DSS for $\mathrm{TMF}_{\Gamma}$. For compactifications of $\mathsf{M}_{\Gamma}$, we need to go to the log étale site of $\overline{\mathsf{M}}_{\mathrm{Ell}}$ as in [HL16]. This is due to the structure map $\overline{\mathsf{M}}_{\Gamma}\to\overline{\mathsf{M}}_{\mathrm{Ell}}$ generally not being étale, but only log étale. In particular, replacing the small étale site of a Deligne–Mumford stack $\mathsf{X}$ with its log étale site, for a chosen log structure on $\mathsf{X}$, all of the results of this article follow mutatis mutandis. This leads to a synthetic $\mathbf{E}_{\infty}$-ring
@@
that, likewise, lifts $\mathrm{Tmf}_{\Gamma}$ and whose signature is the DSS for $\mathrm{Tmf}_{\Gamma}$.
@@
In other words, $\nu$ preserves the limit $\mathcal{O}^{\mathrm{top}}(\mathsf{M}_{\mathrm{Ell}})$ defining $\mathrm{TMF}$. While the synthetic spectrum $\mathrm{SMF}$ is therefore not new, we will continue to denote it by $\mathrm{SMF}$ rather than $\operatorname{\nu}\mathrm{TMF}$, in order to emphasise its construction and the fact that its signature naturally gives the DSS for $\mathrm{TMF}$.
@@
By Pr. 2.14, the cofibre of the map $\operatorname{\nu}\mathrm{Tmf}\to\mathrm{Smf}$ is the $(-1)$-truncation $\tau_{\leq-1}\mathrm{Smf}$. Since $\mathrm{Smf}$ is $(-1)$-connective, this cofibre is therefore the desuspension of the discrete object $\pi_{-1}^{\heartsuit}(\mathrm{Smf})$, considered as an object of $\mathrm{Syn}_{\mathrm{MU}}^{\heartsuit}\simeq\mathrm{Comod}_{\mathrm{MU}_{*}\mathrm{MU}}$. The last claim follows from [Pst23, Th.4.18], which says that if $S$ is a synthetic spectrum, then we have an isomorphism of graded abelian groups
@@
Classically, we have an equivalence of $\mathbf{E}_{\infty}$-rings $\mathrm{Tmf}[\Delta^{-24}]\simeq\mathrm{TMF}$. Proving a synthetic version of such a statement is more subtle, because the synthetic global sections functor does not preserve all colimits (see Rmks. 3.6 and 3.7). Nevertheless, we can show that it does preserve the specific filtered colimits used to invert elements of filtration zero on the global sections of $\mathcal{O}^{\mathrm{syn}}$. This in fact holds for more general even-periodic refinements in the place of $\overline{\mathsf{M}}_{\mathrm{Ell}}$.
@@
Let $(\mathsf{X},\mathcal{O}^{\mathrm{top}}_{\mathsf{X}})$ be a tame even-periodic refinement such that the underlying stack $\mathsf{X}$ is noetherian, and let $f\in\operatorname{\pi}_{2d,0}\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$ be an element detected on the 00-line of the SDSS for $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}$. Write $\overline{f}\in\mathrm{H}^{0}(\mathsf{X},\omega^{\otimes d}_{\mathsf{X}})$ for its image under the edge homomorphism, and write $D(\overline{f})$ for the nonvanishing locus of $\overline{f}$. Then the natural map of synthetic $\mathbf{E}_{\infty}$-$\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$-algebras
$\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})[f^{-1}]\xrightarrow{\cong}\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(D(\overline{f}))$
is an equivalence.
In the proof below, we will identify the signature of $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$ with its DSS via Cor. 2.12. We will moreover abuse notation, and refer to the element in $\operatorname{\pi}_{2d}\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(\mathsf{X})$ obtained by $\tau$-inverting $f\in\operatorname{\pi}_{2d,0}\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$ by the same name.
@@
It suffices to show the natural map above becomes an equivalence after $\tau$-inversion and tensoring with $C\tau$. In the former case, this follows from [MM15, Th.4.14] (this is the only moment where we use the noetherian hypothesis), as this says that the global sections functor $\Gamma(\mathsf{X},-)\colon\mathrm{QCoh}(\mathsf{X},\mathcal{O}_{\mathsf{X}}^{\mathrm{top}})\to\mathrm{Mod}_{\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(\mathsf{X})}(\mathrm{Sp})$ preserves colimits. In particular, we obtain the chain of equivalences
$\mathcal{O}_{\mathsf{X}}^{\mathrm{top}}(\mathsf{X})[f^{-1}]=\Gamma(\mathsf{X},\,\mathcal{O}^{\mathrm{top}}_{\mathsf{X}})[f^{-1}]\cong\Gamma(\mathsf{X},\,\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}[\overline{f}{}^{-1}])\cong\Gamma(D(\overline{f}),\,\mathcal{O}^{\mathrm{top}}_{\mathsf{X}})=\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(D(\overline{f})).$
It suffices to check that $\mathcal{O}_{\mathsf{X}}^{\mathrm{syn}}(\mathsf{X})[f^{-1}]\otimes C\tau\to\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(D(\overline{f}))\otimes C\tau$ is an equivalence, and the cellularity of $\mathrm{Syn}_{\mathrm{MU}}$ allows us to check this on bigraded homotopy groups. As tensoring with $C\tau$ preserves colimits and filtered colimits commute with bigraded homotopy groups, we calculate (using Cor. 2.12) the left-hand side to be the $\mathrm{E}_{2}$-page of the DSS for $\mathcal{O}_{\mathsf{X}}^{\mathrm{top}}(\mathsf{X})$ with $f$ inverted. This agrees with the $\mathrm{E}_{2}$-page of the DSS for $\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}$ at $D(\overline{f})$. Indeed, sheaf cohomology commutes with filtered colimits in the abelian $1$-category of quasi-coherent sheaves (see [Sta, 0GQV]). As $f$ is of filtration zero, it comes from a map in the abelian $1$-category of quasi-coherent sheaves. This $\mathrm{E}_{2}$-page in turn agrees with the homotopy groups of $\mathcal{O}^{\mathrm{syn}}(D(\overline{f}))\otimes C\tau$, and we see that the map induces the natural isomorphism on bigraded homotopy groups. ∎
Let ${\Delta}^{24}\in\operatorname{\pi}_{576,0}\mathrm{Smf}$ denote the unique element that $\tau$-inverts to the usual periodicity element $\Delta^{24}\in\operatorname{\pi}_{576}\mathrm{Tmf}$. Applying the above to $\Delta^{24}$, we obtain a synthetic lift of the usual relationship between $\mathrm{Tmf}$ and $\mathrm{TMF}$.
@@
This follows from Pr. 4.16 and the fact that $\mathsf{M}_{\mathrm{Ell}}$ is exactly the nonvanishing locus of $\overline{\Delta}^{24}$ inside $\overline{\mathsf{M}}_{\mathrm{Ell}}$. ∎
Note that the proof of this statement does not depend on Cor. 4.13, but only uses that $\mathrm{SMF}$ is a synthetic lift of $\mathrm{TMF}$, and that its signature is the DSS for $\mathrm{TMF}$.
We can also compare Df. 4.8 with another possible definition of $\mathrm{Smf}$, one originally proposed by the second author. Recall that $\mathrm{tmf}=\tau_{\geq 0}\mathrm{Tmf}$, and that $c_{4}\in\operatorname{\pi}_{8}\mathrm{tmf}$ is the uniquely defined $\overline{\kappa}$-torsion class detecting the normalised Eisenstein series of weight $4$.
@@
is clearly a pullback, so we are left to identify this square with the square in the statement of the proposition. By Cor. 4.17 and Cor. 4.13, we see that $\mathrm{Smf}[{\Delta}^{-24}]\cong\mathrm{SMF}\cong\operatorname{\nu}\mathrm{TMF}$, respectively. It suffices to now show that the composite morphism of synthetic $\mathbf{E}_{\infty}$-rings

### v2

@@ 3.3   Comparing the DSS and the ANSS
    Using the notation of (3.12), we have an isomorphism of trigraded $\mathbf{Z}[\tau]$-modules [^8] As usual, we set $\omega^{n/2}=0$ for odd $n$.
@@
Using part (2) of Pr. 3.20, we see that assumption (b) is equivalent to the analogous vanishing condition for the $\mathrm{E}_{2}$-page, so (b) certainly implies (a). To prove the converse, suppose that the group $\mathrm{H}^{s}(\mathsf{Y},\omega_{\mathsf{Y}}^{\otimes t})$ does not vanish for certain $s,t$, where $s>n$. By part (2) of Pr. 3.20, this group appears on the $\mathrm{E}_{2}$-page of the SDSS in tridegree $(2t-s,\,s,\,s)$. For degree reasons it cannot be the target of a differential, as differentials only hit elements that are $\tau$-divisible on the $\mathrm{E}_{2}$-page, and this tridegree is not $\tau$-divisible. If this tridegree contains a permanent cycle, then this gives a nonzero contribution to stem $(2t-s,\,s)$, showing that $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$ is not $(-n)$-connective since $s>n$. If, on the other hand, all elements in this tridegree support differentials, then let us pick an element $x$ that supports a $d_{r}^{\mathrm{syn}}$-differential for some $r\geq 2$. Then $d_{r}^{\mathrm{syn}}(x)=\tau^{r-1}\cdot y$ for some class $y$ in tridegree $(2t-s-1,\,s+r,\,s+r)$. By $\tau$-linearity of the differential $d_{r}^{\mathrm{syn}}$, it follows that $y$ is a permanent cycle and gives a nonzero contribution to stem $(2t-s-1,\,s+r)$. Since $s+r>n$, it follows that $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})$ is not $(-n)$-connective. ∎
@@
This identification of spectral sequences, and other similar statements, have appeared elsewhere in the literature:
@@ 4.1   Homotopy fixed-point spectral sequences
that admits a canonical even-periodic refinement given by $\operatorname{Spec}(E)/G$; see [MM15, Pr.2.15]. It is well known that the homotopy fixed-point spectral sequence (HFPSS) for $E$ can be naturally identified with the DSS for the global sections of this even-periodic refinement, which we write as $(\mathsf{X},\mathcal{O}^{\mathrm{top}}_{\mathsf{X}})$ for a moment. Indeed, the cobar resolution computing the homotopy fixed-points $E^{hG}$ takes the form of the cosimplicial object
@@
Let $G$ be a finite group and let $E\in\mathrm{Fun}(BG,\mathrm{CAlg}(\mathrm{Sp}))$ whose underlying ring spectrum is even-periodic and Landweber exact. Then, using the notation above, there is a natural equivalence of cosimplicial $\mathbf{E}_{\infty}$-rings $E^{\bullet}\cong\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(\operatorname{Spec}(\pi_{0}E)^{\bullet})$. In particular, the HFPSS for $E^{hG}$ and the DSS for $\mathsf{X}$ are isomorphic, and the signature spectral sequence of the synthetic $\mathbf{E}_{\infty}$-ring $(\nu E)^{hG}$ is the HFPSS for $E$.
@@
Let $G$ be a finite group and let $E\in\mathrm{Fun}(BG,\mathrm{CAlg}(\mathrm{Sp}))$ whose underlying ring spectrum is even-periodic and Landweber exact. If, for every field-valued point $x$ of $\operatorname{Spec}\pi_{0}E$, the subgroup of $G$ stabilising $x$ acts faithfully on the formal group associated to $x$, then the canonical map
@@
is an equivalence of synthetic $\mathbf{E}_{\infty}$-rings. In particular, the signature spectral sequence of $(\nu E)^{hG}$ is the ANSS of $E^{hG}$.
@@
Remark 4.4.
In both Cor. 4.2 and Th. 4.3, it is not important that $E$ is even-periodic and Landweber exact, only that it is a complex-orientable spectrum.
@@
Example 4.6.
As the $C_{2}$-acting on $\mathrm{KU}$ is a subgroup of the automorphism group of $\widehat{\mathbf{G}}_{m}$, it follows that $G$ always acts faithfully on geometric points of $\operatorname{Spec}\mathbf{Z}$. In particular, it follows that the ANSS for $\mathrm{KO}$ is isomorphic to the $C_{2}$-homotopy fixed-point spectral sequence of $\mathrm{KU}$, a well-known fact. The equivalence of synthetic $\mathbf{E}_{\infty}$-rings $\nu\mathrm{KO}\to(\nu\mathrm{KU})^{hC_{2}}$ was also shown empirically in [CD24, Pr.4.25].
@@
In general, regardless of whether or not $\operatorname{Spec}(\pi_{0}E)/G\to\mathsf{M}_{\mathrm{FG}}$ is affine, the signature spectral sequence of the synthetic spectrum $\nu(E)^{hG}$ is the HFPSS of $E$ by Cor. 4.2. This will not generally coincide with the ANSS of $E^{hG}$, making $\nu(E)^{hG}$ often preferable over $\nu(E^{hG})$ as a synthetic lift of $E^{hG}$. The unit of the synthetic $\mathbf{E}_{\infty}$-ring $\nu(E)^{hG}$ still provides a map of multiplicative spectral sequences
@@
suitable for detection arguments, for example. In this sense, the HFPSS for $E$ is a modified ANSS à la [CD24, Df.2.6]. From this point of view, the ANSS of $E^{hG}$ has no advantage over the HFPSS.
@@
In fact, the corresponding homotopy fixed-point spectral sequence does not coincide with the Adams–Novikov spectral sequence of $\mathrm{KU}^{hC_{2}}\cong\mathrm{KU}^{B{C_{2}}_{+}}$. Indeed, the latter is complex-oriented as a $\mathrm{KU}$-module, and hence its ANSS is concentrated in filtration zero. On the other hand, the cohomology groups $\mathrm{H}^{s}(C_{2},\pi_{*}\mathrm{KU})$ are nontrivial for positive values of $s$, hence the homotopy fixed-point spectral sequence is nontrivial in positive filtrations.
Since the ANSS of $\mathrm{KU}^{B{C_{2}}_{+}}$ collapses on the zero-line, the ANSS provides no computational leverage; computing the $\mathrm{E}_{2}$-page already requires knowing $\pi_{*}(\mathrm{KU}^{B{C_{2}}_{+}})$. The HFPSS, however, has an $\mathrm{E}_{2}$-page that is algebraically computable using only $\pi_{*}\mathrm{KU}$ as input.
@@ 4.2   Synthetic modular forms
By Pr. 2.14, the synthetic $\mathbf{E}_{\infty}$-rings $\mathrm{Smf}$ and $\mathrm{SMF}$ are synthetic lifts of $\mathrm{Tmf}$ and $\mathrm{TMF}$, and by Cor. 2.12, their signature spectral sequences are the DSS for $\mathrm{Tmf}$ and $\mathrm{TMF}$.
@@
We will ignore the case of $\mathrm{tmf}$, as there is no clear way to define this object from the perspective used here. The first main problem is that the stack $\mathsf{M}_{\mathrm{cubic}}$ associated to $\mathrm{tmf}$ is not Deligne–Mumford, as the automorphism group of the cuspidal cubic given by $y^{2}=x^{3}$ is not an étale group scheme. The second problem is that the map $\mathsf{M}_{\mathrm{cubic}}\to\mathsf{M}_{\mathrm{FG}}$ is not flat. To see this, we notice that if it were flat, then $\mathrm{MU}_{\ast}\mathrm{tmf}$ would be flat over $\pi_{\ast}\mathrm{MU}$, which we can see fails once we know that
@@
In a similar vein, $\operatorname{\nu}\mathrm{ko}$ is a preferred $\mathrm{MU}$-synthetic model for $\mathrm{ko}$ as the map $\mathsf{M}_{\mathrm{quad}}\to\mathsf{M}_{\mathrm{FG}}$ (where $\mathsf{M}_{\mathrm{quad}}$ is the stack associated to $\mathrm{ko}$; see [DFHH14, §9]) is also affine. In this case, the first two authors previously showed [CD24, Pr.4.25] by direct computation that there are equivalences
@@
which further confirm this suspicion that $\operatorname{\nu}\mathrm{ko}$ is the preferred synthetic lift of $\mathrm{ko}$; here $\tau_{\geq}^{\uparrow}$ is the vertical $t$-structure of [CD24, Th.3.2]. We repeat the warning from [CD24, Warn.4.26] that these equivalences do not hold in $\mathrm{Syn}_{\mathbf{F}_{2}}$.
@@
Variant 4.13.
It is also possible to alter Df. 4.10 to accommodate moduli stacks of elliptic curves with level structures. Indeed, for a congruence subgroup $\Gamma\subseteq\operatorname{SL}_{2}(\mathbf{Z})$, write $\mathsf{M}_{\Gamma}$ for the moduli stack of elliptic curves with $\Gamma$-level structure. The map $\mathsf{M}_{\Gamma}\to\mathsf{M}_{\mathrm{Ell}}$ is étale, and we define
@@
This is a synthetic $\mathbf{E}_{\infty}$-ring that lifts $\mathrm{TMF}_{\Gamma}$ and whose signature implements the DSS for $\mathrm{TMF}_{\Gamma}$. For compactifications of $\mathsf{M}_{\Gamma}$, we need to go to the log étale site of $\overline{\mathsf{M}}_{\mathrm{Ell}}$ as in [HL16]. This is due to the structure map $\overline{\mathsf{M}}_{\Gamma}\to\overline{\mathsf{M}}_{\mathrm{Ell}}$ generally not being étale, but only log étale. In particular, replacing the small étale site of a Deligne–Mumford stack $\mathsf{X}$ with its log étale site, for a chosen log structure on $\mathsf{X}$, all of the results of this article follow mutatis mutandis. This leads to a synthetic $\mathbf{E}_{\infty}$-ring
@@
that, likewise, lifts $\mathrm{Tmf}_{\Gamma}$ and whose signature induces the DSS for $\mathrm{Tmf}_{\Gamma}$.
@@
In other words, $\nu$ preserves the limit $\mathcal{O}^{\mathrm{top}}(\mathsf{M}_{\mathrm{Ell}})$ defining $\mathrm{TMF}$. While the synthetic spectrum $\mathrm{SMF}$ is therefore not new, we will continue to denote it by $\mathrm{SMF}$ rather than $\operatorname{\nu}\mathrm{TMF}$, in order to emphasise its construction and the fact that its signature implments the DSS for $\mathrm{TMF}$.
@@
By Pr. 2.14, the cofibre of the map $\operatorname{\nu}\mathrm{Tmf}\to\mathrm{Smf}$ is the $(-1)$-truncation $\tau_{\leq-1}\mathrm{Smf}$. Since $\mathrm{Smf}$ is $(-1)$-connective, this cofibre is therefore the desuspension of the discrete object $\pi_{-1}^{\heartsuit}(\mathrm{Smf})$, considered as an object of $\mathrm{Syn}_{\mathrm{MU}}^{\heartsuit}\cong\mathrm{Comod}_{\mathrm{MU}_{*}\mathrm{MU}}$. The last claim follows from [Pst23, Th.4.18], which says that if $S$ is a synthetic spectrum, then we have an isomorphism of graded abelian groups
@@
Classically, we have an equivalence of $\mathbf{E}_{\infty}$-rings $\mathrm{Tmf}[\Delta^{-24}]\cong\mathrm{TMF}$. Proving a synthetic version of such a statement is more subtle, because the synthetic global sections functor does not preserve all colimits (see Rmks. 3.6 and 3.7). Nevertheless, we can show that it does preserve the specific filtered colimits used to invert elements of filtration zero on the global sections of $\mathcal{O}^{\mathrm{syn}}$. This in fact holds for more general even-periodic refinements in the place of $\overline{\mathsf{M}}_{\mathrm{Ell}}$.
@@
Let $(\mathsf{X},\mathcal{O}^{\mathrm{top}}_{\mathsf{X}})$ be a tame even-periodic refinement.
  1. (1)
[…]
    $\left(\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})/\tau\right)[f^{-1}]\xrightarrow{\cong}\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(D(f))/\tau$
    is an isomorphism.
  2. (2)
[…]
    $\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(\mathsf{X})[g^{-1}]\xrightarrow{\cong}\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(D(\overline{g}))$
    is an isomorphism.
@@
For (1), we check that the map induces an isomorphism on bigraded homotopy groups. As filtered colimits commute with bigraded homotopy groups, we calculate (using Cor. 2.12) the left-hand side to be the $\mathrm{E}_{2}$-page of the DSS for $\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(\mathsf{X})$ with $f$ inverted. This agrees with the $\mathrm{E}_{2}$-page of the DSS for $\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}$ at $D(f)$. Indeed, sheaf cohomology commutes with filtered colimits in the abelian $1$-category of quasi-coherent sheaves; see [Sta, Tag 0GQV]. As $f$ is of filtration zero, it comes from a map in the abelian $1$-category of quasi-coherent sheaves. This $\mathrm{E}_{2}$-page in turn agrees with the homotopy groups of $C\tau\otimes\mathcal{O}^{\mathrm{syn}}_{\mathsf{X}}(D(f))$, and we see that the map induces the natural isomorphism on bigraded homotopy groups.
For (2), it suffices to check that the map is an isomorphism after tensoring with $C\tau$ and after $\tau$-inversion. The former of the two follows from (1) for $f=\overline{g}$, because tensoring with $C\tau$ preserves colimits. The latter of the two follows from from [MM15, Th.4.14]; note that this result requires $\mathsf{X}$ to be noetherian and the map $\mathsf{X}\to\mathsf{M}_{\mathrm{FG}}$ to be tame. Indeed, this result says that the global sections functor
$\Gamma(\mathsf{X},-)\colon\mathrm{QCoh}(\mathsf{X},\mathcal{O}_{\mathsf{X}}^{\mathrm{top}})\to\mathrm{Mod}_{\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}(\mathsf{X})}(\mathrm{Sp})$
preserves colimits; in particular, we obtain the chain of natural isomorphisms
$\Gamma(\mathsf{X},\,\mathcal{O}^{\mathrm{top}}_{\mathsf{X}})[f^{-1}]\cong\Gamma(\mathsf{X},\,\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}[\overline{f}{}^{-1}])\cong\Gamma(D(\overline{f}),\,\mathcal{O}^{\mathrm{top}}_{\mathsf{X}}).\qed$
Let ${\Delta}^{24}\in\operatorname{\pi}_{576,0}\mathrm{Smf}$ denote the unique element that $\tau$-inverts to the usual periodicity element $\Delta^{24}\in\operatorname{\pi}_{576}\mathrm{Tmf}$; see [CDvN24, Lm.8.1] for this computation. Applying the above to $\Delta^{24}$, we obtain a synthetic lift of the usual relationship between $\mathrm{Tmf}$ and $\mathrm{TMF}$.
@@
This follows from Pr. 4.18 and the fact that $\mathsf{M}_{\mathrm{Ell}}$ is exactly the nonvanishing locus of $\overline{\Delta}^{24}$ inside $\overline{\mathsf{M}}_{\mathrm{Ell}}$. ∎
Note that the proof of this statement does not depend on Cor. 4.15, but only uses that $\mathrm{SMF}$ is a synthetic lift of $\mathrm{TMF}$, and that its signature implments the DSS for $\mathrm{TMF}$.
We can also compare Df. 4.10 with another possible definition of $\mathrm{Smf}$, one originally proposed by the second author. Recall that $\mathrm{tmf}=\tau_{\geq 0}\mathrm{Tmf}$ and that $c_{4}\in\operatorname{\pi}_{8}\mathrm{tmf}$ is the uniquely defined $\overline{\kappa}$-torsion class detecting the normalised Eisenstein series of weight $4$.
@@
is a pullback. Indeed, by Pr. 4.18, we can identify this square with $\mathcal{O}^{\mathrm{syn}}(D(f))$ for $f=1,\Delta,c_{4}$, and $c_{4}\Delta$, respectively, reading from left-to-right then top-to-bottom. The fact that $\overline{\mathsf{M}}_{\mathrm{Ell}}=D(\Delta)\cup D(c_{4})$ and the fact that $\mathcal{O}^{\mathrm{syn}}$ is a Zariski sheaf then yields the above pullback of synthetic $\mathbf{E}_{\infty}$-rings. We are left to identify this square with the square in the statement of the proposition. By Cor. 4.19 and Cor. 4.15, we see that $\mathrm{Smf}[{\Delta}^{-24}]\cong\mathrm{SMF}\cong\operatorname{\nu}\mathrm{TMF}$, respectively. It suffices to now show that the composite morphism of synthetic $\mathbf{E}_{\infty}$-rings